How much water would a container hold in gallons if the tank is 6 foot by 2 foot by 2 foot? Thank you.
Found 2 solutions by mananth, solver91311:
Answer by mananth(16946)

Volume = L*W*H
V= 6*2*2= 24 cu.feet
..
1cu.ft = 7.48 gallons
24 cu.feet = 7.48*24
=179.52 gallons

Answer by solver91311(24713)



The number of cubic feet in a rectangular solid container is given by the length times the width times the height. Each cubic foot is approximately 7.5 gallons.


John

My calculator said it, I believe it, that settles it